# Asymptotic stability equals exponential stability—while you twist your eyes

### v2

@@ 4   Input-to-State Stability
Leveraging intuition from linear systems theory, input-to-state stability was derived as a stability notion invariant under coordinate transformations (homeomorphisms that fix 00) [ref:sontagbasicconcepts2008input]. In particular, a system (4.1) is said to be Input-to-State Stable (ISS) when there are $\beta\in\mathcal{KL}$ and $\gamma\in\mathcal{K}_{\infty}$ such that
@@
Inspired by the above, it seems plausible that $\dot{x}=-x+d$ fulfils the role of “canonical ISS system” on $\mathbb{R}^{n}$. For instance, recall that the solution satisfies $\varphi^{t}(x_{0},d(\cdot))=e^{-t}x_{0}+\int^{t}_{0}e^{-(t-\tau)}d(\tau)\mathrm{d}\tau$ for any $t\geq 0$, $x_{0}\in\mathbb{R}^{n}$ and $d(\cdot)\in\mathscr{D}$. Therefore, $\|\varphi^{t}(x_{0},d(\cdot))\|\leq\max\{2|e^{-t}|\|x_{0}\|+2\int^{\infty}_{0}|e^{-(t-\tau)}|\mathrm{d}\tau\|d\|_{\infty}\}$. In that sense, $\dot{x}=-x+d$ is canonically ISS due to the simple comparison functions we can employ. Moreover, set $x\mapsto S(x):=\frac{1}{2}\|x\|_{2}^{2}$, it follows that $\langle\partial_{x}S(x),-x+d\rangle\leq\tfrac{1}{2}(\|d\|_{2}^{2}-\|x\|_{2}^{2})$ for all $x,d\in\mathbb{R}^{n}$ and thus
@@
Here, we focus on homotopies through coordinate transformations of $x$ and $d$, for otherwise we could “remove” the disturbance, e.g., consider $f(x,\theta d)$ with $\theta\in[0,1]$. Now, to exemplify why the existence of a homotopy from any ISS system, through coordinate transformation, to $\dot{x}=-x+d$ is too strong, consider the following.
@@
The way to interpret the upcoming proposition is as follows. Given a solution under (4.1), assumed to be ISS. This solution is understood as fixed data and now someone gradually applies a change of coordinates to both $x$ and $d$. It turns out that by doing so, one can always transform this solution data into data satisfying (4.3), thus, into a system like $\dot{x}=-x+d$.
@@
First, by [ref:grune1999asymptotic, Thm. 3] and Proposition 3.2, if (4.1) is ISS, we can always find a continuous path of coordinate transformations of $x$, transforming the solutions to (4.1) into solutions of a system that is ISES with normalized constants and admits a representation of the form (4.1).
@@
On the other hand, for $\|x\|_{2}\leq\alpha(\|d^{\prime}\|_{2})$, (4.7) also holds true. As $d^{\prime}\in D$ was arbitrary, (4.7) holds true for any $d^{\prime}\in D$. Now, let $\widehat{f}(\cdot,v):=f(\cdot,R^{-1}(v))$ and see that if the goal is to get $L_{\widehat{f}(\cdot,v)}V(x)\leq-V(x)+\|v\|_{2}^{2}$, then we could select $\mathbb{R}^{m}\ni z\mapsto R(z):=(\widetilde{\alpha}(\|z\|_{2}))^{1/2}(z/\|z\|_{2})$, with $R(0):=0$, which is $C^{\infty}$ on $\mathbb{R}^{m}\setminus\{0\}$ and continuous at 00 as $\widetilde{\alpha}\in\mathcal{K}_{\infty}$. For $R$ to be orientation preserving, we can again introduce a map $\rho$, e.g., $z\mapsto(\widetilde{\alpha}(\|z\|_{2}))^{1/2}\rho\left(z/\|z\|_{2}\right).$
@@ 5   A view from optimal transport
Suppose there is a smooth Lyapunov function $V$, asserting that $0\in\mathbb{R}^{n}$ is GAS under some ODE $\dot{x}=X(x)$. To refine our understanding beyond Theorem 3.2 and towards a similar result for vector fields, we note that our desire relates to optimal transport (OT). Let $\mu_{0},\mu_{1}\in\mathscr{P}(\mathbb{R}^{n})$ be Borel probability measures on $\mathbb{R}^{n}$. Now suppose that $\mathrm{d}\mu_{0}=\kappa_{0}e^{-V}\mathrm{d}\lambda^{n}$ and $\mathrm{d}\mu_{1}=\kappa_{1}e^{-V_{q}}\mathrm{d}\lambda^{n}$, for $\kappa_{0},\kappa_{1}$ normalization constants. Can we transport $\mu_{0}$ to $\mu_{1}$ along sufficiently regular measures, preserving unimodality? Suppose the answer is yes, then we can construct a path of densities $[0,1]\ni s\mapsto f(x;s)$, with $f(x;0)=e^{-V(x)}$ and $f(x;1)=e^{-V_{q}(x)}$. Now, if $f$ is sufficiently regular, we can conclude that $\nabla\log f(x;s)$ renders 00 GAS, for each fixed $s$. It is particularly interesting that the standard Gaussian measure results in the canonical ODE $\dot{x}=-x$. In what follows we touch upon this viewpoint.
@@
Now consider the density $\mathbb{R}^{n}\ni x\mapsto\rho(x;s):=\kappa(s)^{-1}e^{-\frac{1}{2}\langle x,\Sigma(s)^{-1}x\rangle}$ for $\kappa(s)^{2}:=(2\pi)^{n}\mathrm{det}(\Sigma(s))$ and set $V(\cdot;s):=-\log\rho(\cdot;s)$. Let $\Sigma_{1}=I_{n}$, then it follows that the family of ODEs $\dot{x}=-\nabla V(x;s)$, comprises a homotopy from $\dot{x}=-\nabla V(x;0)$ to $\dot{x}=-x$, through vector fields such that 00 is GAS (e.g., consider the Lyapunov function $x\mapsto\widetilde{V}(x;s):=\frac{1}{2}\langle x,\Sigma(s)^{-1}x\rangle$). $\circ$
@@ 6   Conclusion and future work
We have provided a step towards addressing Conley’s converse question in some generality, yet, many open problems remain. Although directly working with flows has benefits, e.g., see [ref:aguiar2023universal], the main open problem is the extension to vector fields and generic attractors. Several other questions are as follows.
@@
[…]
Open problem 3: prove or disprove that Proposition 3.2 holds for all semiflows that render 00 GAS (not just the ones generated by vector fields).
@@ Appendix A Continuation and the Conley index
In the context of semiflows, Conley’s index theory can be set up as follows [ref:rybakowski1987homotopy, Ch. 1]. Given a semiflow $\varphi$ on $M^{n}$, then, $S\subseteq M^{n}$ is said to be an isolated (positively) invariant set when there is a compact set $K\subseteq M^{n}$, called an isolating neighbourhood, such that $S=\mathrm{I}(K,\varphi):=\{p\in K:\varphi^{t}(p)\in K\,\forall t\geq 0\}\subseteq\mathrm{int}(K)$. Now, a pair of compact sets $(N,L)\subset M^{n}\times M^{n}$ is said to be an index pair for $S$ when (i) $S=\mathrm{I}(\mathrm{cl}(N\setminus L),\varphi)$ and $N\setminus L$ is a neighbourhood of $S$; (ii) $L$ is positively invariant in $N$; and (iii) $L$ is an exit set for $N$. Then, the Conley index of $S$ is the homotopy type of the pointed space $(N/L,[L])$. This index is independent of the choice of index pair. Now, if some $N$ can be chosen to be an isolating neighbourhood throughout a homotopy $[0,1]\ni s\mapsto\varphi(\cdot;s)$, then, the Conley index is preserved along that homotopy and we speak of continuation (of the index). The existence of the homotopy implies continuation, but to what extent does an equivalent Conley index imply the existence of an index preserving homotopy? This is the starting point of this note and this is why we speak of a “converse question”.

### v3

@@ 4  Input-to-State Stability
Leveraging intuition from linear systems theory, input-to-state stability was derived as a stability notion invariant under coordinate transformations (homeomorphisms that fix 0) [ref:sontagbasicconcepts2008input]. In particular, a system (4.1) is said to be Input-to-State Stable (ISS) when there are $\beta\in\mathcal{KL}$ and $\gamma\in\mathcal{K}_{\infty}$ such that
@@
Inspired by the above, it seems plausible that $\dot{x}=-x+d$ fulfils the role of “canonical ISS system” on $\mathbb{R}^{n}$. For instance, recall that the solution satisfies $\varphi^{t}(x_{0},d(\cdot))=e^{-t}x_{0}+\int^{t}_{0}e^{-(t-\tau)}d(\tau)\mathrm{d}\tau$ for any $t\geq 0$, $x_{0}\in\mathbb{R}^{n}$ and $d(\cdot)\in\mathscr{D}$. Therefore, $\|\varphi^{t}(x_{0},d(\cdot))\|\leq\max\{2|e^{-t}|\|x_{0}\|,2\|d\|_{\infty}\}$. In that sense, $\dot{x}=-x+d$ is canonically ISS due to the simple comparison functions we can employ. Moreover, set $x\mapsto S(x):=\frac{1}{2}\|x\|_{2}^{2}$, it follows that $\langle\partial_{x}S(x),-x+d\rangle\leq\tfrac{1}{2}(\|d\|_{2}^{2}-\|x\|_{2}^{2})$ for all $x,d\in\mathbb{R}^{n}$ and thus
@@
Here, we focus on homotopies through coordinate transformations of $x$ and $d$, for otherwise we could “remove” the disturbance, e.g., consider $0,1]\ni\theta\mapsto f(x,\theta d)$. Now, to exemplify why the existence of a homotopy from any ISS system, through coordinate transformation, to $\dot{x}=-x+d$ is too strong, consider the following example. Example 4.1 implies in particular that a naïve generalization of Section 3 to ISS is impossible.
@@
The way to interpret the upcoming proposition is as follows. Given a solution under (4.1), assumed to be ISS. This solution is understood as fixed data and now someone gradually applies a change of coordinates to both $x$ and $d$. It turns out that by doing so, one can always transform this solution data into data satisfying (4.3), thus, into a system like $\dot{x}=-x+d$. Indeed, it follows directly that ISS is preserved throughout, e.g., see [ref:kellett2015input] for more on ISS and coordinate transformations. However, comparing again to Section 3, we cannot readily say more beyond paths of coordinate transformations, e.g., first one should topologize $\mathscr{D}$ to study continuity of $d(\cdot)\mapsto\varphi^{t}(x_{0},d(\cdot))$. In this work we refrain from completing this study as it does not add to the central message while it requires significant technical machinery, however, without doing so we cannot formally discuss homotopies of solutions [^10] This is also why “and ISS equals finite energy” is not part of our title cf. [ref:grune1999asymptotic]..
@@
First, by [ref:grune1999asymptotic, Thm. 3] and Proposition 3.2, if (4.1) is ISS, we can always find a continuous path of coordinate transformations of $x$, transforming the solutions to (4.1) into solutions of a system that is ISES with normalized constants, while still admiting a representation of the form (4.1).
@@
On the other hand, for $\|x\|_{2}\leq\alpha(\|d^{\prime}\|_{2})$, (4.7) also holds true. As $d^{\prime}\in D$ was arbitrary, (4.7) holds true for any $d^{\prime}\in D$. Now, let $\widehat{f}(\cdot,v):=f(\cdot,R^{-1}(v))$ and see that if, towards (4.3), the goal is to get $L_{\widehat{f}(\cdot,v)}V(x)\leq-V(x)+\|v\|_{2}^{2}$, then we could select $\mathbb{R}^{m}\ni z\mapsto R(z):=(\widetilde{\alpha}(\|z\|_{2}))^{1/2}(z/\|z\|_{2})$, with $R(0):=0$. This map is $C^{\infty}$ on $\mathbb{R}^{m}\setminus\{0\}$ and continuous at 0 as $\widetilde{\alpha}\in\mathcal{K}_{\infty}$. For $R$ to be orientation preserving, we can again introduce a map $\rho:\mathbb{S}^{m-1}\to\mathbb{S}^{m-1}$ of appropriate degree, like in Step (ii) of the proof of Proposition 3.2, e.g., $z\mapsto(\widetilde{\alpha}(\|z\|_{2}))^{1/2}\rho\left(z/\|z\|_{2}\right).$
@@ 5  A view from optimal transport
Suppose there is a smooth Lyapunov function $V$, asserting that $0\in\mathbb{R}^{n}$ is GAS under some ODE $\dot{x}=X(x)$. To refine our understanding beyond Theorem 3.2 and towards a similar result for vector fields, we note that our desire relates to optimal transport (OT). Let $\mu_{0},\mu_{1}\in\mathscr{P}(\mathbb{R}^{n})$ be Borel probability measures on $\mathbb{R}^{n}$. Now suppose that $\mathrm{d}\mu_{0}=\kappa_{0}e^{-V}\mathrm{d}\lambda^{n}$ and $\mathrm{d}\mu_{1}=\kappa_{1}e^{-V_{q}}\mathrm{d}\lambda^{n}$, for $\kappa_{0},\kappa_{1}$ normalization constants. Can we transport $\mu_{0}$ to $\mu_{1}$ along sufficiently regular measures, preserving unimodality? Suppose the answer is yes, then we can construct a path of densities $0,1]\ni s\mapsto f(x;s)$, with $f(x;0)=e^{-V(x)}$ and $f(x;1)=e^{-V_{q}(x)}$. Now, if $f$ is sufficiently regular, we can conclude that $\nabla\log f(x;s)$ renders 0 GAS, for each fixed $s$. It is particularly interesting that the standard Gaussian measure results in the canonical ODE $\dot{x}=-x$. In what follows we touch upon this viewpoint.
@@
Now consider the density $\mathbb{R}^{n}\ni x\mapsto\rho(x;s):=\kappa(s)^{-1}e^{-\frac{1}{2}\langle x,\Sigma(s)^{-1}x\rangle}$ for $\kappa(s)^{2}:=(2\pi)^{n}\mathrm{det}(\Sigma(s))$ and set $V(\cdot;s):=-\log\rho(\cdot;s)$. Let $\Sigma_{1}=I_{n}$, then it follows that the family of ODEs $\dot{x}=-\nabla V(x;s)$, comprises a homotopy from $\dot{x}=-\nabla V(x;0)$ to $\dot{x}=-x$, through vector fields such that 0 is GAS (e.g., consider the Lyapunov function $x\mapsto\widetilde{V}(x;s):=\frac{1}{2}\langle x,\Sigma(s)^{-1}x\rangle$).  $\circ$
@@ 6  Conclusion and future work
We have provided a step towards better understanding Conley’s converse question in some generality, yet, many open problems remain. Although directly working with flows has benefits, e.g., see [ref:aguiar2023universal], the main open problem is the extension to vector fields and generic attractors. Several other questions are as follows.
@@
To construct an example of a semiflow that does not correspond to a vector field, one can appeal to maps with a “negative orientation”, e.g., a smooth vector field $X$ always results in a flow $\varphi(\cdot;X)$ such that the diffeomorphism $\varphi^{t}(\cdot;X)$ is isotopic to the identity [^11] Consider the homotopy $0,1]\ni s\mapsto\varphi^{\tau(1-s)}$ from $\varphi^{\tau}$ to $\varphi^{0}=\mathrm{id}$.. To that end, consider the map $\mathbb{R}\ni x\mapsto f(x):=-\alpha x$, for some $\alpha\in(0,1)$, and the define the mapping torus of $(\mathbb{R},f)$ through $M_{f}:=\{(x,t)\in\mathbb{R}\times[0,1]\}/{\sim}$ for $(x,1)\sim(f(x),0)$. Now, the suspension of $f$ (under the ceiling function $c(x)\equiv 1$) is the semiflow $\varphi(\cdot;f)$ defined through $\varphi^{t}((x,s);f):=(f^{n}(x),s^{\prime})$ where the pair $(n,s^{\prime})$ satisfies $n+s^{\prime}=t+s$ with $0\leq s^{\prime}\leq 1$ [ref:brin2002introduction, Sec. 1.11]. It follows that $\{0\}\times[0,1]/{\sim}$ is an attractor.  $\circ$
Open problem 3: prove or disprove that Proposition 3.2 holds for all semiflows that render 0 GAS (not just the ones generated by vector fields).
@@ Appendix A Continuation and the Conley index
In the context of semiflows, Conley’s index theory can be set up as follows [ref:rybakowski1987homotopy, Ch. 1]. Given a semiflow $\varphi$ on $M^{n}$, then, $S\subseteq M^{n}$ is said to be an isolated (positively) invariant set when there is a compact set $K\subseteq M^{n}$, called an isolating neighbourhood, such that $S=\mathrm{I}(K,\varphi):=\{p\in K:\varphi^{t}(p)\in K\,\forall t\geq 0\}\subseteq\mathrm{int}(K)$. Now, a pair of compact sets $(N,L)\subset M^{n}\times M^{n}$ is said to be an index pair for $S$ when (i) $S=\mathrm{I}(\mathrm{cl}(N\setminus L),\varphi)$ and $N\setminus L$ is a neighbourhood of $S$; (ii) $L$ is positively invariant in $N$; and (iii) $L$ is an exit set for $N$. Then, the Conley index of $S$ is the homotopy type of the pointed space $(N/L,[L])$. This index is independent of the choice of index pair. Now, if some $N$ can be chosen to be an isolating neighbourhood throughout a homotopy $0,1]\ni s\mapsto\varphi(\cdot;s)$, then, the Conley index is preserved along that homotopy and we speak of continuation (of the index). The existence of the homotopy implies continuation, but to what extent does an equivalent Conley index imply the existence of an index preserving homotopy? This is the starting point of this note and this is why we speak of a “converse question”.